# Foldable SuperNets: Scalable Merging of Transformers with Different Initializations and Tasks

### v1

@@ Abstract
Many recent methods aim to merge neural networks (NNs) with identical architectures trained on different tasks to obtain a single multi-task model. Most existing works tackle the simpler setup of merging NNs initialized from a common pre-trained network, where simple heuristics like weight averaging work well. This work targets a more challenging goal: merging large transformers trained on different tasks from distinct initializations. First, we demonstrate that traditional merging methods fail catastrophically in this setup. To overcome this challenge, we propose Foldable SuperNet Merge (FS-Merge), a method that optimizes a SuperNet to fuse the original models using a feature reconstruction loss. FS-Merge is simple, data-efficient, and capable of merging models of varying widths. We test FS-Merge against existing methods, including knowledge distillation, on MLPs and transformers across various settings, sizes, tasks, and modalities. FS-Merge consistently outperforms them, achieving SOTA results, particularly in limited data scenarios [^1] Code and models are available at https://github.com/idankinderman/fs_merge.
@@ 1 Introduction
Practitioners frequently train identical neural architectures for various tasks and share these models online, while the original training data is often unavailable due to privacy, proprietary, or other concerns. This led to an increased interest in the field of model merging (Akhlaghi & Sukhov, 2018; Wortsman et al., 2022; Goddard et al., 2024), which aims to combine the weights, and sometimes the features, of several models into a single new model (Figure 1). This approach could allow the merging of multiple single-task models into a single multi-task model (Matena & Raffel, 2022; Ilharco et al., 2023), eliminating the need to store and run multiple models or ensembles (Ganaie et al., 2022).
@@
Most existing merging methods have a strong restriction: they assume that the models were initialized from the same pre-trained model and subsequently fine-tuned. This encourages the models to stay aligned (Ainsworth et al., 2023) and also to remain closer in the weight space ilharco2023editing, and therefore easier to fuse, for example by simply averaging their weights (wortsman2022model). However, this restricts the ability to merge models that do not share the same initialization. For example, consider the task of merging the weights of two unrelated models from an online repository (e.g., Hugging Face or GitHub). These models were likely not fine-tuned from the same initial model, making most existing merging techniques inapplicable.
To overcome this, a few recent studies explored merging differently initialized models using alignment-based methods (entezari2022the; singh2020model; verma2024merging; ainsworth2023git; stoica2024zipit). However, these approaches use very simple merging rules and hence struggle with larger, complex tasks such as merging transformers (stoica2024zipit).
To illustrate the limitations of traditional and alignment-based merging methods, we pre-trained two Vision Transformers (ViTs) (dosovitskiy2021an) with different initializations on ImageNet-1k (deng2009imagenet), then fine-tuned each on separate tasks: Cars (krause20133d) and CIFAR10 (krizhevsky2009learning). Attempts to merge these models into a single multi-task ViT using various methods (Table 1) were severely unsuccessful. Traditional approaches like weight averaging, SLERP (shoemake1985animating), and RegMean (jin2023dataless), as well as the alignment merging method designed for transformers, ‘Opt’ (imfeld2023transformer), all resulted in a merged model with performance comparable to that of a random guess. Moreover, these methods show a significant accuracy gap compared to the model ensembling (ganaie2022ensemble), a method that averages the model outputs. Note that the ensemble is not a valid merging method, as it uses the original models directly. These traditional merging methods rely on simple local merging rules, ensuring computational efficiency. However, they proved inadequate for our complex setting due to their simplicity. This pattern persisted across all other settings, tasks, and modalities we tested with transformers (see Section 3.2).
@@
Table 1: Merging a pair of ViT-B-16, fine-tuned on Cars and CIFAR10, using 100 original training images and 800 augmented images from each dataset. The test accuracy is averaged on both tasks.
Method Accuracy
Ensemble 89.27
Random guess 5.25
Average 5.56
SLERP 4.80
RegMean 6.58
Opt 6.32
Distillation 75.81
FS-Merge (Ours) 84.52
@@
These results indicate that merging large transformers from diverse initializations demands stronger and more resource-intensive techniques, such as Knowledge Distillation (KD) (ba2014deep; hinton2015distilling). In the multi-task setting, KD refers to a single model learning to replicate the outputs of multiple models (tan2018multilingual; vongkulbhisal2019unifying). And indeed, KD significantly outperforms the previous methods (Table 1), but still exhibits a large performance gap compared to the ensemble. Although KD shows promise, it often requires access to the large parts of the original training dataset and labels (clark2019bam; liu2020adaptive) to achieve high accuracy, which may be problematic due to privacy or proprietary concerns. Moreover, it does not explicitly utilize the original models’ weights, potentially overlooking valuable information.
@@
Our approach. In this work, we address the challenging scenario of merging differently initialized single-task transformer (vaswani2017attention) into a unified multitask model of equal size in a data-efficient manner. We propose Foldable SuperNet Merge (FS-Merge), a method that optimizes a SuperNet to fuse the weights of the original networks. This fusion minimizes local or global feature reconstruction loss. Post-optimization, the weights of the original networks are folded into a single merged model (Figure 2). Importantly, our formulation offers greater generality than simple rule-based alignment techniques (stoica2024zipit; imfeld2023transformer). Moreover, by leveraging the original weights, FS-Merge outperforms KD, particularly in data-limited regimes, significantly reducing the gap with the ensemble (Table 1), and even surpassing it in some cases.
FS-Merge is a simple and data-efficient approach that, like other methods (jin2023dataless; ainsworth2023git), requires only an unlabeled fraction of the original training data. Notably, it produces a merged model that maintains the same inference speed and memory usage as those from traditional methods. We demonstrate its effectiveness by achieving SOTA results across multiple scenarios, architectures, model sizes, datasets, and modalities. While this work focuses on transformers, FS-Merge’s versatility allows it to be readily extended to other architectures such as RNNs (sherstinsky2020fundamentals) and CNNs (he2016deep). This adaptability sets FS-Merge apart from alignment-based methods, which are often designed for specific architectures.
@@ 2 Method
For simplicity, we outline the merging problem for two models $A$ and $B$ with identical widths, each trained on distinct tasks with unique initializations and separate classification heads (Figure  1a). We also have an unlabeled subset of the training set from each task, $D^{A}$ and $D^{B}$. Our goal is to develop a new model with the same architecture, that minimizes the losses for tasks $A$ and $B$. Note that the classification heads are not merged, meaning we retain a separate classification head for each task (Figure 1b).
In this section, we introduce FS-Merge, our proposed approach. We begin by explaining how to merge MLPs, a relatively easier task, and then proceed to tackle the more complex challenge of merging transformer models.
@@ 2.1 Warmup: merging Multi-Layer Perceptrons
Figure 2: The FS-Merge for MLP. (a) Local FS-Merge: $M_{l}$ and $U_{l}$ are optimized separately for each layer $l$ to reconstruct $f_{l}^{A}||f_{l}^{B}$. (b) Global FS-Merge: composing the $M$ and $U$ matrices for each one of the $L$ layers of the MLP to reconstruct $f_{L}^{A}||f_{L}^{B}$. In both versions, after optimization, the Foldable SuperNet is folded to create the merged model. Red represents the SuperNet weights, blue represents the original frozen weights, gray represents features, and green represents the activation function.
We first consider the $l$-th layer in the Multi-Layer Perceptron (MLP) model $A$. The features at this layer, denoted by $f_{l}^{A}\in\mathbb{R}^{d_{l}}$, can be expressed as follows:
@@
$\tilde{f}_{l}(z_{l}^{A},\,z_{l}^{B})={\color[rgb]{1,0,0}U_{l}}\sigma({\color[rgb]{1,0,0}M_{l}}(z_{l}^{A}\,||\,z_{l}^{B}))\,,$ (2)
The input is a concatenation of the original pre-activation features $z_{l}^{A}\,||\,z_{l}^{B}\in\mathbb{R}^{2\cdot d_{l}}$. ${\color[rgb]{1,0,0}M_{l}}\in\mathbb{R}^{d_{l}\times 2\cdot d_{l}}$ (“Merge”) is used to merge the original features into a lower-dimensional space (from $2\cdot d_{l}$ to $d_{l}$), and ${\color[rgb]{1,0,0}U_{l}}\in\mathbb{R}^{2\cdot d_{l}\times d_{l}}$ (“Unmerge”) is used to approximately reverse the merge operation, attempting to reconstruct the original post-activation features ${f}_{l}^{A}\,||\,{f}_{l}^{B}\in\mathbb{R}^{2\cdot d_{l}}$ with $\tilde{f}_{l}\in\mathbb{R}^{2\cdot d_{l}}$. Another way to express $\tilde{f}_{l}$ is by $\tilde{f}_{l}^{A}\,||\,\tilde{f}_{l}^{B}$.
For optimization, $D^{A}\cup D^{B}$ is used to extract the models features $f_{l}^{A},$ and $f_{l}^{B}$. Then, $M_{l}$ and $U_{l}$ are optimized separately for each layer $l$, on the following reconstruction optimization problem:
@@
$M_{1}^{*},U_{1}^{*},...M_{L}^{*},U_{L}^{*}=\underset{M_{1},U_{1},...,M_{L},U_{L}}{\text{argmin}}\,\mathbb{E}_{x\sim D}\,\left\|f_{L}^{A}||f_{L}^{B}-\tilde{f}_{L}(x)\right\|_{2}^{2}\,,$ (4)
@@
$W_{l}^{*}=M_{l}^{*}\begin{pmatrix}W_{l}^{A}&0\\ 0&W_{l}^{B}\end{pmatrix}U_{l-1}^{*},\ b_{l}^{*}=M_{l}^{*}\begin{pmatrix}b_{l}^{A}\\ b_{l}^{B}\end{pmatrix},\ U_{0}=\begin{pmatrix}I\\ I\end{pmatrix}\,.$ (5)
Intuitively, this folding operation creates a merged model which “under the hood” reconstructs the original features from the previous layer (using $U_{l-1}$), applies the original weights, and then merges those features again ($M_{l}$), all with the same complexity as each of the original models.
Initialization. The Foldable SuperNet can be initialized in various ways (full details in Appendix G.1). In the simplest approach, the $M$ and $U$matrices may be initialized randomly (“random”). We found that the best approach is to initialize so that only the weights of the first model are selected (“first”). This means that, if we fold the weights at that point, we obtain $W^{*}_{l}=W^{A}_{l}$ This is achieved by initializing $M_{l}$ and $U_{l}$ for each layer $l$ as follows:
$M_{l}=\begin{pmatrix}I&0\end{pmatrix}\,,\,U_{l}=\begin{pmatrix}I\\ I\\ \end{pmatrix}.$ (6)
Relation with ZipIt (stoica2024zipit). This “folding” operation was proposed by ZipIt (stoica2024zipit), which is closely related to our work, and inspired it. This method also merges models from various initializations and tasks, targeting MLPs and CNNs. In the MLP context, ZipIt represents a special case of FS-Merge, where $M$ is chosen to average highly correlated pairs in a hard-coded way, and $U=2M^{\top}$. For other layers, such as skip connections and normalizations, ZipIt employs various heuristics that cannot be easily extended to transformers and may harm performance. For more details, see Appendix A.2.
Extension to multiple models. This methodology can easily be extended to any number of models. It is also capable of merging models of varying widths into any target width dimension, provided they have the same number of layers.
@@ 2.2 Merging transformers
Figure 3: The Foldable SuperNet for merging two attention blocks ($A$ and $B$). Only the red components are trained. After training, we fold the Foldable SuperNet to create the merged block.
@@
Merging transformers (vaswani2017attention) is much more challenging than merging MLPs, due to their larger scale and more complicated structure. Creating a naively Foldable SuperNet as described in the MLP Section 2.1 is not feasible for transformers because it fails to account for their skip connections, layer normalization (ba2016layer), and multi-head attention.
@@
To tackle these issues, we develop a new Foldable SuperNet architecture, and train it using the global objective Eq. 4 (see Appendix B.2 for more details). We found that the global objective is essential in the case of transformers, as training the Foldable SuperNet locally (Eq. 3) significantly reduces the accuracy of the resulting merged model (in Appendix H.1 we suggest a few reasons for this). In this section, we explain how to build a Foldable SuperNet for the attention block (Figure  3), as it is the most complex component of the transformer. For the pre-processing block and the MLP block, please refer to Appendix B.1.
A review of the attention block. Consider the transformer trained on task $A$. The $l$-th attention block gets the features $X_{l}^{A}\in\mathbb{R}^{T\times d}$ as input, when $T$ is the sequence length and $d$ is the embedding size, and applies layer norm with parameters $\gamma^{A}_{l},\beta^{A}_{l}\in\mathbb{R}^{d}$ (Eq. 7 left). Then, it calculates the queries, keys, and values for $H$ heads. An example of query creation is shown in Eq. 7 right.
$\bar{X}_{l}^{A}=\text{LN}_{\gamma^{A}_{l},\beta^{B}_{l}}(X_{l}^{A})\,\,\,;\,\,\,Q^{A}_{l}=\bar{X}_{l}^{A}W_{l}^{Q,A}=(Q_{1,l}^{A},...,Q_{H,l}^{A})\,.$ (7)
Where $Q_{i,l}^{A}\in\mathbb{R}^{T\times\frac{d}{H}}$ is the $i$-th head queries. Similarly, the keys and values are created using $W_{l}^{K,A},\,W_{l}^{V,A}$ respectively. Following this, the transformer executes multi-head attention, utilizes $W_{l}^{O,A}$, and uses a skip connection, which creates the output $Y_{l}^{A}\in\mathbb{R}^{T\times d}$
$Y_{l}^{A}=X_{l}^{A}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{A}{K_{i,l}^{A}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{A},...]W_{l}^{O,A}\,.$ (8)
Foldable SuperNet for attention blocks. Suppose there are two transformers trained on distinct tasks $A$ and $B$. Our goal is to define a Foldable SuperNet that will use the original weights and will allow us to merge the attention blocks of these two transformers after training. Intuitively, this Foldable SuperNet will reconstruct the original features of models $A$ and $B$, create queries keys and values, learn how to merge them, and apply multi-head attention. The optimized parameters at this stage are highlighted in red. The notation $\tilde{X}$ represents a reconstruction attempt of our method, and $X^{*}$ represents the merged features or parameters.
The $l$-th attention block gets merged features from the previous layer as input, $X_{l}^{*}\in\mathbb{R}^{T\times d}$. Then, the Foldable SuperNet learns the parameters of the layer norm of this block ${\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}\in\mathbb{R}^{d}$, along with a matrix ${\color[rgb]{1,0,0}U_{l}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the features from the merged input
$\tilde{X}_{l}^{A}||\tilde{X}_{l}^{B}=\text{LN}_{{\color[rgb]{1,0,0}\gamma^{*}_{l},\beta^{*}_{l}}}(X_{l}^{*}){\color[rgb]{1,0,0}U_{l}}\,.$ (9)
Observe that the layer norm in each block is also being optimized, similar to what has been proposed in other merging works (jordan2023repair). This does not hurt the efficiency of our method due to the small number of learnable parameters in the layer norm. Then, the Foldable SuperNet calculates the queries, keys, and values for $A$ and $B$. These components, drawn from all heads and models, are then concatenated and merged using learnable matrices ${\color[rgb]{1,0,0}M_{l}^{Q},M_{l}^{K},M_{l}^{V}}\in\mathbb{R}^{2\cdot d\times d}$. For example, the queries are merged as follows:
$\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}=(\tilde{X_{l}}^{A}W_{l}^{Q,A}||\tilde{X_{l}}^{B}W_{l}^{Q,B})\in\mathbb{R}^{T\times 2\cdot d}\,,\,Q_{l}^{*}=(\tilde{Q}^{A}_{l}||\tilde{Q}^{B}_{l}){\color[rgb]{1,0,0}M_{l}^{Q}}=(Q_{1,l}^{*},...,Q_{H,l}^{*})\,.$ (10)
@@
Then, they are divided into $H$ heads. Similarly, the keys and values are compressed using the matrices ${\color[rgb]{1,0,0}M_{l}^{K},M_{l}^{V}}$ respectively. Following this, the Foldable SuperNet executes multi-head attention in the original width of $d$, thanks to the compression of queries, keys and values; uses ${\color[rgb]{1,0,0}U_{l}^{O}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original multi-head attention output; utilizes $W_{l}^{O,A}\,,\,W_{l}^{O,B}$; uses ${\color[rgb]{1,0,0}M_{l}^{O}}\in\mathbb{R}^{2\cdot d\times d}$ to compress it once more; and applies a skip connection. Together, we get:
$Y_{l}^{*}=X_{l}^{*}+\text{Concat}_{i}[...,\text{softmax}\left(\frac{Q_{i,l}^{*}{K_{i,l}^{*}}^{\top}}{\sqrt{d}}\right)V_{i,l}^{*},...]\ {\color[rgb]{1,0,0}U_{l}^{O}}\begin{pmatrix}W_{l}^{O,A}&0\\ 0&W_{l}^{O,B}\end{pmatrix}{\color[rgb]{1,0,0}M_{l}^{O}}\,.$ (11)
This process results in $Y_{l}^{*}\in\mathbb{R}^{T\times d}$, serving as the input for the subsequent MLP block at layer $l$. It is important to note that this Foldable SuperNet was designed so the skip connection and layer norm are applied to the compressed features, and that for any $M$ and $U$, there is a linear layer that comes before or after it. This arrangement will allow us to fold this structure into the merged model after training (see Appendix B.3 for more details).
@@
Parameterizing $M$ and $U$. Utilizing full-rank $M_{l}\in\mathbb{R}^{d_{l}\times n\cdot d_{l}}$ and $U_{l}\in\mathbb{R}^{n\cdot d_{l}\times d_{l}}$ for merging $n$ models introduces a number of parameters that increase quadratically with the layer width $d_{l}$. In large models like transformers, this leads to a very high demand for resources and hinder the optimization process. To mitigate this, we adopt a parameterization strategy akin to LoRA (hu2022lora), using a sum of a low-rank matrix and a concatenation of diagonal matrices. For instance, in $M_{l}$:

### v2

@@ Abstract
Recent methods aim to merge neural networks (NNs) with identical architectures trained on different tasks into a single multi-task model. While most works focus on the simpler setup of merging NNs initialized from a common pre-trained network, we target the harder problem of merging large transformers trained on different tasks from distinct initializations. We show that traditional merging methods fail catastrophically in this setup, while Knowledge Distillation (KD) achieves much better results, though at a higher cost. However, KD is data-inefficient, as it does not exploit the original models’ weights. To solve this, we introduce “Foldable SuperNet Merge” (FS-Merge), which trains a SuperNet containing the original models (with frozen weights) using a feature reconstruction objective. After training, the SuperNet is folded back to the size of a single original model. FS-Merge is simple, data-efficient, has a computational cost comparable to KD, and is proven to have superior expressiveness compared to traditional merging methods on MLP models. It achieves SOTA results when tested on MLPs and transformers across various sizes, tasks, modalities, and distribution shifts, especially in low-data scenarios [^1] Code is available at https://github.com/idankinderman/fs_merge..
@@ 1 Introduction
Practitioners frequently train identical neural architectures for various tasks and share these models online, while the original training data is often unavailable due to privacy, proprietary, or other concerns. This led to an increased interest in the field of model merging (Akhlaghi & Sukhov, 2018; Goddard et al., 2024), which aims to combine the weights, and sometimes the features, of several models into a single new model (Figure 1). This approach could allow the merging of multiple single-task models into a single multi-task model (Matena & Raffel, 2022; Ilharco et al., 2023), eliminating the need to store and run multiple models via Mixture-of-Experts Shazeer et al. (2017) or ensembles (Ganaie et al., 2022).
@@
Most existing merging methods have a strong restriction: they assume that the models were initialized from the same pre-trained model and subsequently fine-tuned. This encourages the models to stay aligned (Ainsworth et al., 2023) and also to remain closer in the weight space Ilharco et al. (2023), and therefore easier to fuse, for example by simply averaging their weights (Wortsman et al., 2022). However, this restricts the ability to merge models that do not share the same initialization. For example, consider the task of merging the weights of two unrelated models from an online repository (e.g., Hugging Face or GitHub). These models were likely not fine-tuned from the same initial model, making most existing merging techniques inapplicable.
@@
Table 1: Merging a pair of ViT-B-16, fine-tuned on Cars and CIFAR10, using 100 original images and 800 augmented images from each dataset. The test accuracy is averaged on both tasks.
Method Accuracy Method type
Ensemble 89.27 -
Random guess 5.25 -
Average 5.56 Averaging-based
SLERP 4.80 Averaging-based
RegMean 6.58 Averaging-based
Opt 6.32 Alignment-based
Distillation 75.81 Training-based
FS-Merge (Ours) 85.83 Training-based
To overcome this, several studies have explored merging differently initialized models using alignment-based methods (Entezari et al., 2022; Verma & Elbayad, 2024; Ainsworth et al., 2023; Stoica et al., 2024). This family of methods is more resource-intensive compared to averaging-based methods. However, these approaches still use relatively simple merging rules and thus struggle with more complicated tasks such as merging transformers (Stoica et al., 2024). For a detailed discussion of related work, refer to Appendix A.1.
To illustrate the limitations of these methods, we pre-trained two Vision Transformers (ViTs) (Dosovitskiy et al., 2021) with different initializations on ImageNet-1k (Deng et al., 2009), fine-tuned each on separate tasks (Cars (Krause et al., 2013) and CIFAR10 (Krizhevsky et al., 2009)), and then merged them (Table 1). Traditional approaches like weight averaging, SLERP (Shoemake, 1985), and RegMean (Jin et al., 2023), as well as the alignment merging method designed for transformers, ’Opt’ (Imfeld et al., 2023), all failed in this setting, resulting in a merged model with performance comparable to random guessing. Moreover, these methods show a significant accuracy gap compared to the model ensembling (Ganaie et al., 2022), a method that averages the model outputs. Note that the ensemble is not a valid merging method, as it uses the original models directly. These traditional merging methods rely on simple local merging rules, ensuring computational efficiency. However, they proved inadequate for our complex setting due to their simplicity. This pattern persisted across all other settings, tasks, and modalities we tested with transformers (see Section 3.2).
@@
These results indicate that merging large transformers from diverse initializations demands stronger and more resource-intensive techniques, such as Knowledge Distillation (KD) (Ba & Caruana, 2014; Hinton et al., 2015). In the multi-task setting, KD refers to a single model trained to replicate the outputs of multiple models (Tan et al., 2019; Vongkulbhisal et al., 2019). And indeed, KD significantly outperforms the previous methods (Table 1), but still exhibits a large performance gap compared to the ensemble. Although KD shows promise, it often requires access to large parts of the original training dataset and labels (Clark et al., 2019; Liu et al., 2020) to achieve high accuracy, which may be problematic due to privacy or proprietary concerns. Moreover, it does not explicitly utilize the original models’ weights, potentially overlooking valuable information.
Our approach. To leverage the hidden knowledge in the original weights and further improve accuracy beyond KD, we propose the “Foldable SuperNet Merge (FS-Merge)” method. This approach utilizing a “SuperNet”— a network in which the original models are embedded with frozen parameters. The SuperNet weights are then trained to minimize either local or global feature reconstruction loss. After training, the SuperNet weights are folded to produce a model with the same size as the original models (Figure 2). Notably, this means that the additional computational cost occurs only once, during merging.
@@
FS-Merge is simple, data-efficient, and we prove that it has superior expressivity compared to the baselines (Appendix H). Like other methods (Jin et al., 2023; Ainsworth et al., 2023), it requires only an unlabeled fraction of the original training data. We demonstrate its effectiveness by achieving SOTA results across diverse scenarios, architectures, model sizes, tasks, modalities, and distribution shifts.
We prove that FS-Merge offers greater expressiveness compared to traditional merging methods. Moreover, by leveraging the original weights, FS-Merge outperforms KD, especially in data-limited regimes, significantly reducing the gap with the ensemble (Table 1), and even surpassing it in some cases. While this work focuses on transformers, FS-Merge’s versatility allows it to be extended to other architectures such as RNNs (Sherstinsky, 2020) and CNNs (He et al., 2016). This adaptability sets FS-Merge apart from alignment-based methods, which are often designed for specific architectures.
@@ 2 Method
For simplicity, we outline the merging problem for two models $A$ and $B$ with identical widths, trained on distinct tasks with unique initializations and separate classification heads (Figure 1a). Using unlabeled subsets of each task’s training set, $D^{A}$ and $D^{B}$, our goal is to construct a new model with the same architecture, that minimizes the losses for tasks $A$ and $B$. Be aware that the classification heads are not merged, meaning we retain a separate head for each task (Figure 1b). Note that FS-Merge can be easily extended to merge any number of models or models with varying widths. We denote $D=D^{A}\cup D^{B}$.
In this section, we introduce FS-Merge, our proposed approach, first for merging MLPs and then for the more complex task of merging transformers.
@@ 2.1 Warmup: merging Multi-Layer Perceptrons
Let us consider the $l$-th layer in the Multi-Layer Perceptron (MLP) model $A$. The features at this layer, denoted by $f_{l}^{A}\in\mathbb{R}^{d_{l}}$, can be expressed as follows:
@@
Figure 2: The FS-Merge for MLP. (a) Local FS-Merge: $M_{l}$ and $U_{l}$ are optimized separately for each layer $l$ to reconstruct $f_{l}^{A}||f_{l}^{B}$. (b) Global FS-Merge: composing the $M$ and $U$ matrices for each one of the $L$ layers of the MLP to reconstruct $f_{L}^{A}||f_{L}^{B}$, the features before the classification head. In both versions, after optimization, the Foldable SuperNet is folded to create the merged model. Red represents the SuperNet weights, blue represents the original frozen weights, gray represents features, and green represents the activation function.
@@
$\tilde{f}_{l}(z_{l}^{A},\,z_{l}^{B})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}}\sigma({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}}(z_{l}^{A}\,||\,z_{l}^{B}))\,$, (2)
The input is a concatenation of the original pre-activation features $z_{l}^{A}\,||\,z_{l}^{B}\in\mathbb{R}^{2\cdot d_{l}}$. ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}}\in\mathbb{R}^{d_{l}\times 2\cdot d_{l}}$ (“Merge”) is used to merge the original features into a lower-dimensional space (from $2\cdot d_{l}$ to $d_{l}$), and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}}\in\mathbb{R}^{2\cdot d_{l}\times d_{l}}$ (“Unmerge”) is used to approximately reverse the merge operation, attempting to reconstruct the original post-activation features ${f}_{l}^{A}\,||\,{f}_{l}^{B}\in\mathbb{R}^{2\cdot d_{l}}$ with $\tilde{f}_{l}\in\mathbb{R}^{2\cdot d_{l}}$. Another way to express $\tilde{f}_{l}$ is by $\tilde{f}_{l}^{A}\,||\,\tilde{f}_{l}^{B}$.
For optimization, $D$ is used to extract the models features $f_{l}^{A},$ and $f_{l}^{B}$. Then, $M_{l}$ and $U_{l}$ are optimized separately for each layer $l$, on the following reconstruction optimization problem:
@@
$M_{1}^{*},U_{1}^{*},...,M_{L}^{*},U_{L}^{*}=\\ \underset{M_{1},U_{1},...}{\text{argmin}}\,\mathbb{E}_{x\sim D}\,\left\|f_{L}^{A}||f_{L}^{B}-\tilde{f}_{L}(x)\right\|_{2}^{2}\,$, (4)
@@
Figure 3: The Foldable SuperNet for merging two attention blocks (AA and $B$). Only the red components are trained. $M$ matrices are used to merge the models’ features into merged features ($X^{*}$), and $U$ are applied to the merged features to reconstruct the original ones ($\tilde{X}^{A},\tilde{X}^{B}$). After training, the Foldable SuperNet is folded to produce a merged attention block with the same size as the original blocks.
@@
$W_{l}^{*}=M_{l}^{*}\begin{pmatrix}W_{l}^{A}&0\\ 0&W_{l}^{B}\end{pmatrix}U_{l-1}^{*},\ b_{l}^{*}=M_{l}^{*}\begin{pmatrix}b_{l}^{A}\\ b_{l}^{B}\end{pmatrix}\,$. (5)
Where $U_{0}=\begin{pmatrix}I\\ I\end{pmatrix}$. Intuitively, this folding operation creates a merged model which “under the hood” reconstructs the original features from the previous layer (using $U_{l-1}$), applies the original weights, and then merges those features again ($M_{l}$). This occurs with the same complexity as each of the original models, since $W_{l}^{*}\in\mathbb{R}^{d_{l}\times d_{l-1}}$.
Initialization. The Foldable SuperNet can be initialized in various ways (full details in Appendix G.1). In the simplest approach, the $M$ and $U$ matrices may be initialized randomly. We found that the best approach is to initialize $M$ and $U$ using the “first” initialization method (as described in Eq. 6), where only the weights of the first model are selected. This implies that, if the weights are folded at this point, we obtain $W^{*}_{l}=W^{A}_{l}$. For an ablation study in the matter, see Appendix G.1 and Appendix G.2.
$M_{l}=\begin{pmatrix}I&0\end{pmatrix}\,,\,U_{l}=\begin{pmatrix}I\\ I\\ \end{pmatrix}$ (6)
Relation with ZipIt (Stoica et al., 2024). This “folding” operation was proposed by ZipIt (Stoica et al., 2024), which is closely related to our work, and inspired it. This method also merges models from various initializations and tasks, targeting MLPs and CNNs. In the MLP context, ZipIt represents a special case of FS-Merge, where $M$ is chosen to average highly correlated pairs in a hard-coded way, and $U=2M^{\top}$. For other layers, such as skip connections and normalizations, ZipIt employs various heuristics that cannot be easily extended to transformers and may harm performance. For more details, see Appendix A.2.
@@ 2.2 Merging transformers
Merging transformers (Vaswani et al., 2017) is a much more challenging problem than merging MLPs, due to their larger scale and more complicated structure. Creating a naively Foldable SuperNet as described in the MLP Section 2.1 is not feasible for transformers because it fails to account for their skip connections, layer normalization (Ba et al., 2016), and multi-head attention.
@@
To tackle these issues, we develop a new Foldable SuperNet architecture, and train it using the global objective (Eq. 4). We found that the global objective is essential for merging transformers, as training the SuperNet locally (Eq. 3) significantly reduces the accuracy of the resulting merged model (reasons for this are discussed in Appendix G.3).
Foldable SuperNet for attention blocks. In this section, we briefly describe the Foldable SuperNet for the attention block (Figure 3), as it is the most complex component of the transformer. For full technical details on the SuperNet for this block, as well as for the other transformer blocks (MLP and pre-processing), please refer to Appendix B.1. The optimized parameters at this stage are highlighted in red, with $\tilde{X}$ represents a reconstruction attempt of our method, and $X^{*}$ represents the merged features.
Suppose there are two transformers trained on distinct tasks $A$ and $B$. The Foldable SuperNet for their attention blocks receives as input $X_{l}^{*}\in\mathbb{R}^{T\times d}$, the merged features from the previous SuperNet, where $T$ is the sequence length and $d$ is the embedding size. It then applies a layer norm (which is also learned, as proposed in other merging works (Jordan et al., 2023)), and uses a learned matrix ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original model features $\tilde{X}_{l}^{A}||\tilde{X}_{l}^{B}\in\mathbb{R}^{T\times 2\cdot d}$.
@@
Next, the Foldable SuperNet calculates the queries, keys, and values for $A$ and $B$ using the original frozen weights. The queries of both models are then concatenated and merged using a learnable matrix ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{Q}}\in\mathbb{R}^{2\cdot d\times d}$, and similarly for the keys and values with $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{K},M_{l}^{V}$}. Following this, the Foldable SuperNet performs multi-head attention with the merged queries, keys, and values; uses ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U_{l}^{O}}\in\mathbb{R}^{d\times 2\cdot d}$ to reconstruct the original multi-head attention output; applies the aggregation frozen weights of the original models; uses ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{l}^{O}}\in\mathbb{R}^{2\cdot d\times d}$ to compress it once more; and applies the skip connection. This process results in merged features of dimension $T\times d$, serving as the input for the subsequent Foldable SuperNet of the MLP block at layer $l$.
@@
Parameterizing $M$ and $U$. Utilizing full-rank $M_{l}\in\mathbb{R}^{d_{l}\times n\cdot d_{l}}$ and $U_{l}\in\mathbb{R}^{n\cdot d_{l}\times d_{l}}$ for merging $n$ models introduces a number of parameters that increase quadratically with the layer width $d_{l}$. In large models like transformers, this leads to a very high demand for resources, hinders the optimization process, and cause overfitting. To mitigate this, we adopt a parameterization strategy akin to LoRA (Hu et al., 2022), using a sum of a low-rank matrix and a concatenation of diagonal matrices. For instance, in $M_{l}$: